# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*

### v3

@@ VII Conclusion
This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user aims to minimize the cost function while the adversary tries to maximize it. Additionally, we introduce the RDDPG algorithm, which integrates RDPG with TD3 to improve both robustness and learning efficiency. In order to evaluate the robustness of RDDPG, we implement it on the tracking tasks of the quadrotor. Experimental results show that RDDPG can achieve the optimal control policy and maintain stability in various environments with disturbances. This demonstrates that RDDPG outperforms other DRL methods in terms of robustness.

### v4

@@ 7 Conclusion
This paper proposes robust deterministic policy gradient (RDPG), which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce robust deep deterministic policy gradient (RDDPG), which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.
Appendix A Hyperparameters Used in Training
A.1 Physical parameters of Crazyflies 2.0
Table 1: Physical parameters of the Crazyfile 2.0
Parameter Description Value
m Total mass 0.0270.027[kg]
l Arm length 0.03970.0397[m]
$K_{F}$ Thrust coefficient $3.16\times 10^{-10}$
$K_{M}$ Torque coefficient $7.94\times 10^{-12}$
$I_{xx}$ Principal Moment of Inertia around x axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$I_{yy}$ Principal Moment of Inertia around y axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$I_{zz}$ Principal Moment of Inertia around z axis $2.17\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$K_{D.xy}$ Drag coefficients in $x,y$ axes $9.18\times 10^{-7}$
$K_{D.z}$ Drag coefficients in $z$ axis $10.31\times 10^{-7}$
A.2 Cost function
Table 2: Scaling factors in cost function
Parameter Description Value
$\alpha$ Coefficient for position error $10$
$\beta$ Coefficient for velocity error 11
$\epsilon$ Coefficient for angle error 11
$\lambda$ Living reward 10-10
$\zeta$ Penalty $1000$
A.3 DDPG and TD3
Table 3: Caption
Parameter Description Value
$\alpha_{actor}$ Learning rate for user 0.0001
$\beta_{actor}$ Learning rate for adversary 0.0001
$\alpha_{critic}$ Learning rate for critic 0.001
$|B|$ Batch size 256
$\gamma$ Discounted factor 0.99
$\sigma$ Exploration noise 0.2
A.4 SAC
Table 4: SAC hyper-parameters
Parameter Description Value
$\alpha_{actor}$ Learning rate for user 0.0003
$\alpha_{critic}$ Learning rate for critic 0.003
$|B|$ Batch size 256
$\gamma$ Discounted factor 0.99
$\alpha_{entropy}$ Entropy coefficient 0.99
Appendix B Experiment results
Table 5: The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}$ between the target and the quadrotor. Bold indicates the best performance algorithm
$\pm 5$ wind speed $\pm 10$ wind speed $\pm 15$ wind speed
Traj (1) Traj (2) Traj (1) Traj (2) Traj (1) Traj (2)
[…]
Figure 6: In $\pm 15$ m/s wind scenario
Figure 7: In $\pm 15$ m/s wind scenario
Figure 8: In $\pm 15$ m/s wind scenario
Figure 9: In $\pm 15$ m/s wind scenario
Figure 10: In $\pm 15$ m/s wind scenario